OpenStudy (anonymous):

The compound inequality 8.00 ≤ x < 9.50 represents all values, x, for which college students are paid hourly as teacher assistants. What is another way of writing this compound inequality? x ≤ 8.00 and x < 9.50 x ≤ 8.00 or x < 9.50 x ≥ 8.00 and x < 9.50 x ≥ 8.00 or x < 9.50

OpenStudy (anonymous):

\[a \le x \le b\] means \[a \le x~and~x \le b~or~ x \ge a~and~x \le b\]

OpenStudy (michele_laino):

hint: we can read the original compound inequality as follows: the set of all values \(x\), such that \(x\) is greater or equal than \(8.00\), \(and\) \(x\) is less than \(9.50\)